OpenStudy (anonymous):

Check my work: Find the orthogonal trajectories of the given family of curves: y^2-x^2=c So I use implicit differentiation to get \[\frac{ dy }{ dx } = \frac{ x }{ y }\] Then, I take the negative reciprocal \[\frac{ dy }{ dx } = \frac{ -y }{ x }\] Next, I separate the terms and integrate so \[\int\limits_{}^{} -\frac{ 1 }{ y }dy = \int\limits_{}^{}\frac{ 1 }{ x }dx\] giving \[-lny = lnx +c\]

OpenStudy (anonymous):

exponentiating both sides, Ive got: \[\frac{ 1 }{ y } = xe ^{c}\]. Now I just call e^c C, because its still some constant. So now I've got \[\frac{ 1 }{ y } =cx\] and then just divide by x to get c alone. So i end with \[\frac{ 1 }{ xy } = c\] How does this look? Am I close?

OpenStudy (anonymous):

Nobody knows how to do this?